hi can some please explain this to me.
find the x-intercept of y=x^2-4x-32 by factoring.
Thank you

Answer by Alan3354(69443)

find the x-intercept of y=x^2-4x-32 by factoring.
-----------
Factor x^2-4x-32
Find a pair of factors of 32 that differ by 4
1*32 NG
2*16 NG
4*8 works
At the x-intercepts, y = 0
(x+4)*(x-8) = 0
x = -4
x = 8